OpenStudy (anonymous):

Prove the identity: -tan^2x + sec^2x = 1

OpenStudy (anonymous):

I have no clue where to start.

OpenStudy (anonymous):

\[-\tan ^{2}x + \sec ^{2}x = 1\]

zepdrix (zepdrix):

We have 3 important formulas that relate squares,\[\large \sin^2x+\cos^2x=1\]\[\large \tan^2x+1=\sec^2x\]\[\large \cot^2x+1=\sec^2x\] Can we maybe use one of those identities to help us solve this?

OpenStudy (anonymous):

The second one?

zepdrix (zepdrix):

Yah that seems like a good idea. So what we want to do to "Prove" is to leave one side alone, and try to match it with the other side.

zepdrix (zepdrix):

\[\large -\tan ^{2}x + \color{orangered}{\sec ^{2}x} = 1\]Let's replace this sec^2x using our identity,\[\large \sec^2x=\color{orangered}{\tan^2x+1}\]

OpenStudy (anonymous):

\[-\tan^2x + \tan^2x + 1 = 1\]

OpenStudy (anonymous):

Is that right?

zepdrix (zepdrix):

\[\large -\tan^2x+\color{orangered}{\tan^2x+1}=1\]Yes very good ^^

OpenStudy (anonymous):

Yay! So now what?

zepdrix (zepdrix):

Hmm we have a tan^2, and a negative tan^2.. let's combine them.

OpenStudy (anonymous):

So they cancel each other out, right? You're left with 1 = 1.

zepdrix (zepdrix):

Yesss \c:/ yay team!

OpenStudy (anonymous):

Yay! Thank you so much! You are a life saver. :)